A Complete Bibliography of The Bell System Technical Journal, 1940–1949

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

16 March 2020 Version 1.04

Title word cross-reference

8000-Cycle [LPSW49].

A-C [Bar49]. Abstracts [Ano40a, Ano40b, Ano40c, Ano40d, Ano41a, Ano41b, Ano41d, Ano41c, Ano42a, Ano43a, Ano43b, Ano43c, Ano44a, Ano44b, Ano44c, Ano44d, Ano45a, Ano45b, Ano45c, Ano46a, Ano46b, Ano46c, Ano47a, Ano47b, Ano47c, Ano47d, Ano48a, Ano48b, Ano48c, Ano48d, Ano49a, Ano49b, Ano49c, Ano49d]. Accuracy [Slo40]. Action [BB49]. Active [Pet48]. Admittances [Rob49a, Rob49c]. Advances [Mat40]. Amplification [Hol49]. Amplifier [Bod40, PH49]. Amplifiers [Ben40, For46]. Analogies [Mas41].

Analyses [Sta41]. Analysis [Bud48, Dar40a, Gou40, Loz47, McS44, Ric44, Ric45, Sch48]. angle [Hoy47]. Antennas [FL47, MT47]. Aperiodic [Har41]. Application [GS40, Ric49b]. Applications [DM48, Mas43]. approximate [GS48]. Arbitrarily [BP49]. Arrays [Sch43]. Arsenic [SP40]. Articles [Ano40a, Ano40b, Ano40c, Ano41a, Ano41b, Ano41d, Ano41c, Ano42a, Ano43a, Ano43b, Ano43c, Ano44a, Ano44b, Ano44c, Ano44d, Ano45a, Ano45b, Ano45c, Ano46a, Ano46b, Ano46c, Ano47a, Ano47b, Ano47c, Ano47d, Ano48a, Ano48b, Ano48c, Ano48d, Ano49a, Ano49b, Ano49c, Ano49d]. Aspect [Clo48]. Aspects [RK49, SS44]. associated [Ric49b, Tho47]. Atoms [Dar41]. Attenuation [Bod40, Tho47]. Authors [Ano40a, Ano40b, Ano40c, Ano40d, Ano41a, Ano41b, Ano41d, Ano41c, Ano42a, Ano43a, Ano43b, Ano43c, Ano44a, Ano44b, Ano44c, Ano44d, Ano45a, Ano45b, Ano45c, Ano46a, Ano46b, Ano46c, Ano47a, Ano47b, Ano47c, Ano47d, Ano48a, Ano48b, Ano48c, Ano48d, Ano49a, Ano49b, Ano49c, Ano49d].

Balancing [Van43]. Ballistics [Pie45, Sam45]. Band [FHS49, FB49, GK44, Mum49]. Based [Gou40]. Basis [Boz40]. Beam [Sea48, Ske44]. Behavior [Clo48, Eds45]. Bell [Ano40a, Ano40b, Ano40c, Ano40d, Ano41a, Ano41b, Ano41d, Ano41c, Ano42a, Ano43a, Ano43b, Ano43c, Ano44a, Ano44b, Ano44c, Ano44d, Ano45a, Ano45b, Ano45c, Ano46a, Ano46b, Ano46c, Ano47a, Ano47b, Ano47c, Ano47d, Ano48a, Ano48b, Ano48c, Ano48d, Ano49a, Ano49b, Ano49c, Ano49d]. Below [Ben40]. Bending [Egg45]. bends [Ric48a]. Between [Bod40, BO40, Hol49]. Biased [Ben47]. Book [LeC43]. Branching [LT48]. Broad [Mum49]. Broad-Band [Mum49]. Buried [Sun45, TS49].

C [Bar49]. Cable [BO40, Sun45, TS49, Wad40]. Cables [Gou40, Heb41]. Calcium [SB41]. Carrier [DR49, Dud40, FHS49, Heb41, JP46, LPSW49, Mat40]. Cathode [Fay47]. Cavities [WSK46]. Cavity [KW47b, KW47a]. Centimeter [Eng44, FHH46]. Changes [GK44]. Characteristics [For46, Heb41, JP46]. Circuit [RK49]. Circuited [Gou40]. Circuits [CML41, Cur40, DR49, Fai45, Pet48, Pol40, Sha49b, Wis48]. circular [KW47a, Ric48a]. Classical [Dar43a]. Clock [Mar48]. Coaxial [BO40, Gou40]. Code [Goo47, MP48, Sea48]. Coefficient [Mas40]. Coefficients [Wis48]. Coil [Pet46]. Coils [LG40]. Commercial [Pol40]. Communication [Sha48a, Sha48b, Sha49a]. complex [Hoy47]. Component [NP40]. Compressed [LG40]. Concentrations [SB41]. Conductors [BO40]. conformal [Ric49a]. connected [Ric49b]. Conquest [Sha44]. Constant [Cla46]. Constants [Eng44]. Contact [Bar49, Cur40, Leu46]. Contributions [Ful46]. Contributors [Ano40e, Ano40f, Ano40g, Ano40h, Ano41e, Ano41f, Ano41g, Ano41h, Ano42b, Ano43d, Ano43e, Ano43f, Ano44e, Ano44f, Ano44g, Ano44h, Ano45d, Ano45e, Ano45f, Ano46d, Ano46e, Ano46f, Ano46g, Ano47e, Ano47f, Ano47g, Ano47h, Ano48e, Ano48f, Ano48g, Ano48h, Ano49e, Ano49f, Ano49g, Ano49h]. Control [TH46]. Copper [Blo41, Pad42, Wil41b]. corners [Ric49a]. Corona [Ske40]. Coronaviser [Ske40]. Coupling [Syk44a]. Criterion [Slo49]. Cross [Ben40]. Cross-Modulation [Ben40]. Crossarms [Egg45]. Crosstalk [BO40, Gou40]. Cryptography [Sha45]. Crystal [Bon43a, Bur40, D’H44, Gre44, Mar48, Mas43, SO47]. Crystals [Arm46, Bon43c, BA43, Fai45, MS40, Mas40, Mas47, Syk44a, Wil44]. Culture [Leu46]. Current [TS49]. currents [KW47a]. Curved [Alb49]. Cushioning [Min45]. Cycle [LPSW49]. cylinder [KW47a]. Cylindrical [KW47b].

Decay [Leu46]. Decibel [Joh46]. Deflection [Sea48]. Delay [DR49, Har41, Koc48, Rin48]. Delays [Rap46]. Design [Blo41, Bod40, Bro49, Rid48, Syk44a, Van43, Wil41b]. Destroying [Leu46]. Determination [Bon43a, SB41]. Determining [BA43, Wil44]. Development [SO47]. Deviations [D’H44]. Dialing [Clo48]. Diamond [Pad42]. Dielectric [Eng44]. Dies [Pad42]. differential [GS48, Ric49b]. Diffraction [GS40]. Digests [GS40]. Dimensions [Wer41]. Direct [AL49]. Discernibility [GK44]. Discharge [SCM46]. Distance [Sha44]. Distortion [Rea44, Rin48]. Divided [MS40]. Division [Ben41]. Double [DR41, KC42]. Drawing [Blo41, Pad42, Wil41b]. Drying [Wad40]. Dynamics [Min45, Mac43].

Early [TH46]. Editorial [Ano49i]. Effect [Bla43, Clo48, NP40, Rea44]. Effects [D’H44, Syk44a]. Eight [DR49]. Eight-Kilocycle [DR49]. Electric [Bon43b]. Electrical [Bla41, MS40, Mas41, Wad40]. Electrodes [MS40]. Electromagnetic [LeC43, Sch48]. Electron [GS40, Hol49, Pie45, Pie48a, Rob49a, Sam45, Sea48, Slo49, Mac43]. Electronic [Rob49a]. Electrostatic [BP49]. Elements [BP49]. Elimination [FHS49]. Employing [MS40]. Encountered [Rap46]. End [KW47a]. Engineering [CML41, Wer41]. Entropy [Dar42]. Equalization [DR49]. Equation [Pie43]. Equations [Mas47, GS48, Mac43, Ric41, Ric49b]. Equilizers [Bro49]. Equipping [Wil41b]. Equivalent [Pet48]. Etch [Wil44]. Evaluating [Egg45]. Evolution [Mar48]. Excess [Her49]. Experimental [Hol49, MP48]. Expressed [Mas47]. Expressions [Cla46]. Extension [Hei40].

Fabrication [Gre44]. Facilities [GFF46]. Factors [Eng44]. Fair [SMG40]. Faults [Vaa40]. Feedback [Bla43, Bod40]. Ferromagnetism [Boz40]. Field [BP49, Sch48]. Fields [Pie48b, Sam45]. Filamentary [SPH49]. Fillet [SBP43]. Filter [FHS49, LT48]. Filters [Bur40, D’H44, MS40, Mum48]. Fire [TH46]. Fire-Control [TH46]. First [Mas47]. Fission [Dar40b]. Flat [Mum48]. FM [Zin48]. Focused [Ske44]. Forces [Dar41]. Form [Mas47]. Four [Pet48, Rob49c]. Four-Pole [Rob49c]. Four-Terminal [Pet48]. Free [McS44]. Frequencies [AL49, Heb41, Rob49b]. Frequency [DM48, For46, PH49, Sam45]. Full [Ske40]. functions [Hoy47]. fundamental [Mac43]. Fungi [Leu46]. Future [Buc42].

Gap [GHDR46]. Gas [SCM46].

Gas-Discharge [SCM46]. Generator [FHH46]. Germanium [SPH49]. Ground [Wis48]. guide [Ric49b]. Guides [Alb49, Ric48a, Ric49a, Ric49b].

Harmonic [Sta41]. Having [McS44]. Hearing [SMG40]. Heterodyne [Slo40]. High [Blo41, Fay47, LG40, Pad42, PH49, Sam45, Slo40, WSK46, Mac43]. High-Frequency [PH49, Sam45]. High-Speed [Pad42]. High-Vacuum [Fay47]. History [Bla41]. Holding [Wil41a]. Hole [SPH49]. Holes [Her49]. horn [Ric49b]. Hybrid [Bud48].

Ideal [Ben47]. Idle [Rap46]. II [Wil41b]. Images [Bal40]. Impedance [Bar49, Bla43, Pie43, Pom47]. Imperfections [Wil43]. Index [Har45]. Indicator [CGM40]. Indicial [Mot44]. Inductance [LG40]. Industrial [Fry41]. infinite [Tho47]. Injection [SPH49]. Inspection [DR41, KC42, Wil43]. Instrument [Ske40]. Insulation [Lit41]. Interaction [Hol49]. Interesting [GS40]. Intermediate [For46]. Intermittent [Eds45]. Involved [BB49]. Ionosphere [Dar40a]. Isotropic [McS44]. Issue [Ano40e, Ano40f, Ano40g, Ano40h, Ano41e, Ano41f, Ano41g, Ano41h, Ano42b, Ano43d, Ano43e, Ano43f, Ano44e, Ano44f, Ano44g, Ano44h, Ano45d, Ano45e, Ano45f, Ano46d, Ano46e, Ano46f, Ano46g, Ano47e, Ano47f, Ano47g, Ano47h, Ano48e, Ano48f, Ano48g, Ano48h, Ano49e, Ano49f, Ano49g, Ano49h].

Joints [SBP43]. Junction [Sho49].

Junctions [Sho49].

Kilocycle [DR49].

Laboratory [Leu46]. Layers [Arm46]. Lead [SP40, SB41]. Lengths [Eng44]. Lenses [Koc48]. Level [CGM40]. Lightning [Sun45, TS49]. Limitations [Pie45]. Line [Cla46, Pie43, Ric41]. Linear [Ben44, Gra46, Pet48, Sch43, GS48]. Lines [SW41]. Links [Hei40]. Locations [Vaa40]. Low [Mas40].

Machines [Blo41]. Magnetically [Ske44]. Magnetron [FHH46]. Manufacture [Bur40]. Manufacturing [D’H44, KC42]. Margins [Rea44]. Materials [Bon43b, SE40]. Mathematical [Ric44, Ric45, Sch43, Sha45, Sha48a, Sha48b]. Mathematics [Bon43c, Fry41]. Matrix [Ric48a]. Maximally [Mum48]. Maximally-Flat [Mum48]. Means [Bon43a]. Measurement [Har45, Pom47, Rin48]. Measurements [Wil41a]. Measuring [AL49, Rob49b]. Mechanical [Mas41, Mul48]. Mechanics [Dar43b]. Megacycles [Rob49a]. Memorial [Dar43a]. Metallic [Koc48, SE40]. Metallurgy [SBP43, SS44]. Method [Rob49b]. Methods [Bon43a, JP46, Sch48, Syk44a]. Microwave [Fri48, GFF46, Mum49, Rin48, Rob47, Rob49b, Rob49c, SO47, WSK46]. Microwaves [Bud48, LT48]. Mill [Wil41b]. Mineral [Bon43b]. Mismatches [Pom47]. Modes [McS44, Syk44a]. Modulated [Loz47]. Modulation [Ben40, DF49, Goo47, MP48, Sea48]. Modulator [Fay47]. modulus [Hoy47]. Molybdenum [LG40]. Motion [Syk44a, Mac43]. Mounting [Gre44, Syk44b]. Multichannel [Ben40, MP48]. Multiplex [Ben41]. Musa [Pol40].

n [Sho49]. Nature [Dud40]. Naval [TH46].

Navigation [Rob47]. Network [FHS49]. Networks [Pet48, Van43]. Noise [Ben44, Mum49, Pie48a, Ric44, Ric45, Ric48b]. Non [LT48]. Non-reflecting [LT48]. Normal [MS40, Hoy47]. Note [Ano49i, Pie43, Rid48]. Novel [Cla46]. Nuclear [Dar40b].

Observation [Hol49]. Observations [TS49]. Observing [Ske40]. Obtaining [Rap46]. Open [Gou40]. Operating [Wil41b]. Operation [Blo41, Pol40, Rea44, Wil41b]. Operators [Rap46]. Optical [Bon43a]. Order [Mas47, Ric49b]. Orientation [Bon43a, BA43, Wil44]. Oscillator [Fai45]. Oscillators [DF49, Eds45, PS47, Slo40]. Overload [Ben40]. Oxide [Fay47]. Oxide-Cathode [Fay47].

p [Sho49]. Package [Min45]. Paper [Lit41]. Papers [Ano40d]. Parallel [Rid48, Rob49a]. Parallel-Plane [Rob49a]. Parallel-Tuned [Rid48]. Part [Blo41, Wil41b]. particles [Mac43]. Pass [FHS49]. Passive [Rob49c]. Performance [Bud48, FB49, Har45, JP46, Slo49]. Permalloy [LG40]. Phase [AL49, Bod40, DF49, FHS49, Rob49b, Vaa40, Tho47]. Phenomena [Cur40, Her49]. Physical [BB49, Bon43c, Boz40, Pie45]. Physicist [Dar41]. Piezo [Bon43b]. Piezo-Electric [Bon43b]. Piezoelectric [Fai45, Mas47]. Plane [Rob49a]. Plant [Blo41, Clo48, KC42, Wil41b]. plate [KW47a]. Plated [Gre44]. Plates [Har45, McS44, Syk44b]. plus [Ric48b]. Pole [Rob49c]. Polyrod [MT47]. Potential [Wis48]. Powder [SS44]. Power [Eng44, Vaa40]. Powered [LG40]. Precise [AL49]. Precision [Pom47]. Preservation [Leu46]. Principles [BB49, Syk44b]. Probability [Hoy47]. Problems [Wer41]. Producers [Sta41]. Program [CML41, DR49, FHS49, GK44, LPSW49].

Propagation [Alb49, Cla46]. Properties [BGP47, Bon43c, Ric48b]. Protection [Sun45]. Pulp [Lit41]. Pulse [Fay47, Goo47, Loz47, MP48, Sea48]. Pulsers [Pet46].

Q [WSK46]. Quadrature [NP40]. Quality [LG40, MP48]. Quantitative [SB41, SPH49]. Quantized [Ben48]. Quartz [Bon43a, BA43, Bur40, Gre44, Har45, Mar48, Mas40, Mas43, Syk44a, Syk44b, Wil43, Wil44].

Radar [For46, FL47, GHDR46, GFF46, Mor47, Pet46, Rob47, SO47]. Radars [TH46]. Radial [Ske44]. Radio [DM48, Hei40, Pol40]. Random [Ric44, Ric45, Ric48b]. Rapid [SB41]. Raw [Wil43]. Ray [Arm46]. Rays [BA43]. Reactance [DF49]. Reactive [Bro49]. Reading [AL49]. Receive [SCM46]. Receiver [Mor47, Zin48]. Receivers [Mot44, Rea44, SO47]. Receiving [Pol40]. Rectangular [McS44, Ric48a, Ric49a, Ric49b]. Rectifier [Bar49, Ben44, Ben47]. Rectifiers [SO47]. Reference [CGM40]. reflecting [LT48]. Reflection [Ric49a]. Reflections [Ric48a, Ric49b]. Reflex [PS47]. Regarding [Ano49i]. Related [Har41]. Relations [Bod40]. Relative [Egg45]. Reliability [Wil41a]. Repeater [Fri48]. Repeaters [Rin48]. Requirements [Ben40, CML41]. Research [Fri48, Ful46]. Resistances [Pie48a]. Resistors [BGP47]. Resonance [Bla41]. Resonant [WSK46]. resonator [KW47a]. Resonators [KW47b]. Response [Ben44, Mot44, Zin48]. Results [KW47b, SMG40]. Return [Wis48]. Review [LeC43]. Rings [Bud48]. Rubber [Ful46].

S. [LeC43]. Sampling [DR41, KC42]. Schelkunoff [LeC43]. Second [Mas47, Ric49b]. second-order [Ric49b].

Secrecy [Sha49a]. semi [Tho47]. semi-infinite [Tho47]. Semiconductor [Her49]. Semiconductors [Ano49i, Sho49]. Sensitive [BGP47]. Servo [Gra46]. set [Ric49b]. Shapes [Sta41]. Sharpness [Bal40]. Shift [DM48, DF49]. Short [Gou40]. Short-Circuited [Gou40]. side [KW47a]. Sideband [NP40, Pol40]. Signal [Ben44]. Signals [Ben48, Har41]. Silicon [SO47]. Simulated [Bal40]. Simulating [FHS49]. sine [Ric48b]. Single [DR41, NP40, Pol40]. slope [Tho47]. Small [SB41]. Soil [Leu46]. Solar [Ske40]. Solder [SP40]. Soldered [SBP43]. Solution [Vaa40, GS48]. solutions [Ric41]. Some [Cla46, Ful46, KW47b, RK49, SS44, Sta41]. Source [Mum49]. Spaced [BP49]. Spark [GHDR46]. Specifying [Bon43a]. Spectra [Ben48]. Spectrum [Loz47]. Speech [Dud40]. Speed [Blo41, Pad42, Mac43]. Standard [CGM40]. Start [Rea44]. Start-Stop [Rea44]. State [Har41, Ric41]. Statistical [Dar43b, Ric48b]. Statistics [Dar43a]. Steady [Har41, Ric41]. Stop [Rea44]. Streams [Hol49, Pie48a]. Strength [Egg45]. Studies [Arm46, SPH49]. Study [Leu46, Rap46]. Subjective [Bal40]. Subscribers [Clo48]. Sunlight [Ske40]. Surface [Arm46]. Surfaces [McS44]. Survey [Bon43b]. Switch [SCM46]. Switches [GHDR46]. Switching [Cur40, Sha49b]. Synthesis [Sha49b]. Synthetic [Ful46]. System [AL49, Hei40, LPSW49, MP48, Pol40, SE40, Ano40a, Ano40b, Ano40c, Ano40d, Ano41a, Ano41b, Ano41c, Ano42a, Ano43a, Ano43b, Ano43c, Ano44a, Ano44b, Ano44c, Ano44d, Ano45a, Ano45b, Ano45c, Ano46a, Ano46b, Ano46c, Ano47a, Ano47b, Ano47c, Ano47d, Ano48a, Ano48b, Ano48c, Ano48d, Ano49a, Ano49b, Ano49c, Ano49d]. Systems [Ben41, FHS49, Mul48, Sha49a, Vaa40, Ano41d].

Tables [DR41, Joh46, Tho47]. Targets [Rob47]. TE01 [Alb49]. Technical [Ano40a, Ano40b, Ano40c, Ano40d, Ano41a, Ano41b, Ano41d, Ano41c, Ano42a, Ano43a, Ano43b, Ano43c, Ano44a, Ano44b, Ano44c, Ano44d, Ano45a, Ano45b, Ano45c, Ano46a, Ano46b, Ano46c, Ano47a, Ano47b, Ano47c, Ano47d, Ano48a, Ano48b, Ano48c, Ano48d, Ano49a, Ano49b, Ano49c, Ano49d, GS40]. Technique [Leu46, Wil44]. Techniques [GFF46]. Telegraph [JP46, Mat40, Rea44]. Telegraphy [DM48]. Telephone [Cur40, Heb41, Hei40, Lit41, Mot44, Pol40, SE40, Wad40]. Telephony [Buc42, Goo47, Sha44]. Television [Bal40, SW41]. Temperature [Mas40]. Tensor [Mas47]. Terminal [Pet48, Sha49b, Van43]. Terms [Pie43]. Tertiaries [Gou40]. Test [SB41]. Testing [GFF46, WSK46]. Tests [SMG40]. Their [Bon43a]. Theoretical [McS44]. Theory [Bar49, Gra46, Her49, Sch43, Sha45, Sha48a, Sha48b, Sha49a, Sho49, Ric48a]. Thermally [BGP47]. Thermistors [BGP47]. Three [Vaa40]. Three-Phase [Vaa40]. Time [Ben41, Wil41a]. Tin [SP40]. Tolerances [Wer41]. Toll [Heb41, MP48, Rap46, Sun45]. Transatlantic [Pol40]. Transconductance [Slo49]. transformation [Ric49a]. Transformer [Rid48]. Transient [Her49, Zin48]. Transients [Mul48]. Transistor [BB49, RK49]. Transistors [SPH49, Sho49]. Transmission [AL49, CML41, FB49, Heb41, LPSW49, Mat40, NP40, Pie43, SW41, Ric41]. Transmit [SCM46]. Transmit-Receive [SCM46]. Transoceanic [Buc42]. Transport [Her49]. Transverse [Pie48b]. Traveling [Pie48b]. Traveling-Wave [Pie48b]. Triodes [Rob49c]. Trunk [Clo48, Rap46]. Tube [DF49, Sea48, Ske44, Slo49]. Tubes [BP49, Fay47, For46, Pie48b, Rob49a]. Tuned [Rid48]. Twinning [Wil44]. Two [Hol49, Sha49b, Vaa40, Van43].

Two-Terminal [Van43]. Type [PH49].

Uniform [Cla46]. unit [Tho47]. Units [D’H44, Gre44]. Use [BA43, Wil44]. Used [Sta41]. Uses [BGP47]. Using [KC42].

Vacuum [BP49, Fay47, For46, Ske44]. variate [Hoy47]. Various [JP46]. Vessels [TH46]. Vibration [McS44]. Video [AL49]. Visual [SB41]. Volume [CGM40].

wall [KW47a]. Wave

[Alb49, Eng44, MS40, Pie48b, Sta41, Ric48a, Ric48b, Ric49a, Ric49b]. Wave-Lengths [Eng44]. Waveguide [Bud48, Mum48, Pom47]. Waveguide-Hybrid [Bud48]. Waves [Alb49, FHH46, LeC43, Loz47]. Width [FB49, GK44]. Wiped [SBP43]. Wiping [SP40]. Wire [Blo41, DM48, Lit41, Pad42, Sha44, SW41, Wil41b]. Wood [Leu46]. Wood-Destroying [Leu46]. World [Dar41, SMG40].

X [Arm46, BA43]. X-Ray [Arm46]. X-Rays [BA43].

References

Alsberg:1949:PDR

[AL49] D. A. Alsberg and D. Leed. A precise direct reading phase and transmission measuring system for video frequencies. The Bell System Technical Journal, 28 (2):221–238, April 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                                  tronic). URL http://bstj.bell-
                                                  labs.com/BSTJ/images/Vol28/
                                                  bstj28-2-221.pdf;   http://
                                                  www.alcatel-lucent.com/bstj/

vol28-1949/articles/bstj28-
2-221.pdf.

Albersheim:1949:PTW

[Alb49] W. J. Albersheim. Propagation of TE01 waves in curved wave guides. The Bell System Technical Journal, 28(1):1–32, January 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol28-
1949/articles/bstj28-1-1.pdf.

Anonymous:1940:ATAa

[Ano40a] Anonymous. Abstracts of technical articles from Bell System authors. The Bell System Technical Journal, 19(1):156–158, January 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-1-156.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
1-156.pdf.

Anonymous:1940:ATAb

[Ano40b] Anonymous. Abstracts of technical articles from Bell System authors. The Bell System Technical Journal, 19(3):489–491, July 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-3-489.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
3-489.pdf.

Anonymous:1940:ATAc

[Ano40c] Anonymous. Abstracts of technical articles from Bell System authors. The Bell System Technical Journal, 19(4):647–650, October 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol19/
                                        bstj19-4-647.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol19-1940/articles/bstj19-
                                        4-647.pdf.

Anonymous:1940:ATP

[Ano40d] Anonymous. Abstracts of technical papers by Bell System authors. The Bell System Technical Journal, 19(2):336–337, April 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol19/
                                        bstj19-2-336.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol19-1940/articles/bstj19-
                                        2-336.pdf.

Anonymous:1940:CIa

[Ano40e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 19(1):159–160, January 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol19/
                                        bstj19-1-159.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol19-1940/articles/bstj19-
                                        1-159.pdf.

Anonymous:1940:CIb

[Ano40f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 19(2):338–339, April 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-2-338.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
2-338.pdf.

Anonymous:1940:CIc

[Ano40g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 19(3):492–493, July 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-3-492.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
3-492.pdf.

Anonymous:1940:CId

[Ano40h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 19(4):651, October 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-4-651.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
4-651.pdf.

Anonymous:1941:ATAa

[Ano41a] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 20(1):125–128, January 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-1-125.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
1-125.pdf.

Anonymous:1941:ATAb

[Ano41b] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 20(2):250–252, April 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-2-250.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
2-250.pdf.

Anonymous:1941:ATAd

[Ano41c] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 20(4):439–441, October 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-4-439.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
4-439.pdf.

Anonymous:1941:ATAc

[Ano41d] Anonymous. Abstracts of technical articles by Bell Systems authors. The Bell System Technical Journal, 20(3):359–363, July 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-3-359.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
3-359.pdf.

Anonymous:1941:CIa

[Ano41e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 20(1):129–130, January 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-1-129.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
1-129.pdf.

Anonymous:1941:CIb

[Ano41f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 20(2):253–254, April 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-2-253.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
2-253.pdf.

Anonymous:1941:CIc

[Ano41g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 20(3):364, July 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/

                                        bstj20-3-364.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol20-1941/articles/bstj20-
                                        3-364.pdf.

Anonymous:1941:CId

[Ano41h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 20(4):442, October 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol20/
                                        bstj20-4-442.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol20-1941/articles/bstj20-
                                        4-442.pdf.

Anonymous:1942:ATA

[Ano42a] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 21(1):75–78, June 1942. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol21/
                                        bstj21-1-75.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol21-
                                        1942/articles/bstj21-1-75.pdf.

Anonymous:1942:CI

[Ano42b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 21(1):79, June 1942. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol21/
                                        bstj21-1-79.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol21-
                                        1942/articles/bstj21-1-79.pdf.

Anonymous:1943:ATAa

[Ano43a] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 22(1):136–142, January 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-1-136.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
1-136.pdf.

Anonymous:1943:ATAb

[Ano43b] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 22(2):266–267, July 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-2-266.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
2-266.pdf.

Anonymous:1943:ATAc

[Ano43c] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 22(3):397–401, October 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-3-397.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
3-397.pdf.

Anonymous:1943:CIa

[Ano43d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 22(1):143, January 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol22/
                                        bstj22-1-143.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol22-1943/articles/bstj22-
                                        1-143.pdf.

Anonymous:1943:CIb

[Ano43e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 22(2):268, July 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol22/
                                        bstj22-2-268.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol22-1943/articles/bstj22-
                                        2-268.pdf.

Anonymous:1943:CIc

[Ano43f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 22(3):402, October 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol22/
                                        bstj22-3-402.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol22-1943/articles/bstj22-
                                        3-402.pdf.

Anonymous:1944:ATAa

[Ano44a] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 23(1):130–132, January 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-1-130.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
1-130.pdf.

Anonymous:1944:ATAb

[Ano44b] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 23(2):203–205, April 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-2-203.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
2-203.pdf.

Anonymous:1944:ATAc

[Ano44c] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 23(3):333–335, July 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-3-333.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
3-333.pdf.

Anonymous:1944:ATAd

[Ano44d] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 23(4):458–459, October 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol23/
                                        bstj23-4-458.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol23-1944/articles/bstj23-
                                        4-458.pdf.

Anonymous:1944:CIa

[Ano44e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 23(1):133, January 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol23/
                                        bstj23-1-133.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol23-1944/articles/bstj23-
                                        1-133.pdf.

Anonymous:1944:CIb

[Ano44f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 23(2):206, April 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol23/
                                        bstj23-2-206.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol23-1944/articles/bstj23-
                                        2-206.pdf.

Anonymous:1944:CIc

[Ano44g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 23(3):336, July 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol23/

bstj23-3-336.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
3-336.pdf.

Anonymous:1944:CId

[Ano44h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 23(4):460, October 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-4-460.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
4-460.pdf.

Anonymous:1945:ATAa

[Ano45a] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 24(1):157–158, January 1945. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol24/
bstj24-1-157.pdf;   http://
www.alcatel-lucent.com/bstj/
vol24-1945/articles/bstj24-
1-157.pdf.

Anonymous:1945:ATAb

[Ano45b] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 24(2):301–302, April 1945. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol24/
bstj24-2-301.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol24-1945/articles/bstj24-
                                        2-301.pdf.

Anonymous:1945:ATAc

[Ano45c] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 24(3):462– 466, July/October 1945. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol24/
                                        bstj24-3-462.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol24-1945/articles/bstj24-
                                        3-462.pdf.

Anonymous:1945:CIa

[Ano45d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 24(1):159, January 1945. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol24/
                                        bstj24-1-159.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol24-1945/articles/bstj24-
                                        1-159.pdf.

Anonymous:1945:CIb

[Ano45e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 24(2):303, April 1945. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol24/
                                        bstj24-2-303.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol24-1945/articles/bstj24-
                                        2-303.pdf.

Anonymous:1945:CIc

[Ano45f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 24(3):467, July/October 1945. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol24/
bstj24-3-467.pdf;   http://
www.alcatel-lucent.com/bstj/
vol24-1945/articles/bstj24-
3-467.pdf.

Anonymous:1946:ATAa

[Ano46a] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 25(1):159–164, January 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-1-159.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
1-159.pdf.

Anonymous:1946:ATAb

[Ano46b] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 25(3):532–536, July 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-3-532.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
3-532.pdf.

Anonymous:1946:ATAc

[Ano46c] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 25(4):652–654, October 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol25/
                                        bstj25-4-652.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol25-1946/articles/bstj25-
                                        4-652.pdf.

Anonymous:1946:CIa

[Ano46d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 25(1):165–166, January 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol25/
                                        bstj25-1-165.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol25-1946/articles/bstj25-
                                        1-165.pdf.

Anonymous:1946:CIb

[Ano46e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 25(2):349, April 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol25/
                                        bstj25-2-349.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol25-1946/articles/bstj25-
                                        2-349.pdf.

Anonymous:1946:CIc

[Ano46f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 25(3):537–538, July 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-3-537.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
3-537.pdf.

Anonymous:1946:CId

[Ano46g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 25(4):655–656, October 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-4-655.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
4-655.pdf.

Anonymous:1947:ATAa

[Ano47a] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 26(1):213–216, January 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-1-213.pdf;   http://
www.alcatel-lucent.com/bstj/
vol26-1947/articles/bstj26-
1-213.pdf.

Anonymous:1947:ATAb

[Ano47b] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 26(2):388–393, April 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol26/
                                        bstj26-2-388.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol26-1947/articles/bstj26-
                                        2-388.pdf.

Anonymous:1947:ATAc

[Ano47c] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 26(3):682–690, July 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol26/
                                        bstj26-3-682.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol26-1947/articles/bstj26-
                                        3-682.pdf.

Anonymous:1947:ATAd

[Ano47d] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 26(4):900–903, October 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol26/
                                        bstj26-4-900.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol26-1947/articles/bstj26-
                                        4-900.pdf.

Anonymous:1947:CIa

[Ano47e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 26(1):217–218, January 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol26/
bstj26-1-217.pdf;   http://
www.alcatel-lucent.com/bstj/
vol26-1947/articles/bstj26-
1-217.pdf.

Anonymous:1947:CIb

[Ano47f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 26(2):394, April 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-2-394.pdf;   http://
www.alcatel-lucent.com/bstj/
vol26-1947/articles/bstj26-
2-394.pdf.

Anonymous:1947:CIc

[Ano47g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 26(3):691, July 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-3-691.pdf;   http://
www.alcatel-lucent.com/bstj/
vol26-1947/articles/bstj26-
3-691.pdf.

Anonymous:1947:CId

[Ano47h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 26(4):904, October 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-4-904.pdf;   http://
www.alcatel-lucent.com/bstj/

                                        vol26-1947/articles/bstj26-
                                        4-904.pdf.

Anonymous:1948:ATAa

[Ano48a] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 27(1):175–180, January 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-1-175.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        1-175.pdf.

Anonymous:1948:ATAb

[Ano48b] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 27(2):372–376, April 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-2-372.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        2-372.pdf.

Anonymous:1948:ATAc

[Ano48c] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 27(3):589–590, July 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-3-589.pdf;   http://
                                        www.alcatel-lucent.com/bstj/

vol27-1948/articles/bstj27-
3-589.pdf.

Anonymous:1948:ATAd

[Ano48d] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 27(4):747–751, October 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-4-747.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
4-747.pdf.

Anonymous:1948:CIa

[Ano48e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 27(1):181–182, January 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-1-181.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
1-181.pdf.

Anonymous:1948:CIb

[Ano48f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 27(2):377–378, April 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-2-377.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
2-377.pdf.

Anonymous:1948:CIc

[Ano48g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 27(3):591–592, July 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-3-591.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        3-591.pdf.

Anonymous:1948:CId

[Ano48h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 27(4):752, October 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-4-752.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        4-752.pdf.

Anonymous:1949:ATAa

[Ano49a] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 28(1):157–161, January 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/
                                        bstj28-1-157.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol28-1949/articles/bstj28-
                                        1-157.pdf.

Anonymous:1949:ATAb

[Ano49b] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 28(2):329–331, April 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-2-329.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
2-329.pdf.

Anonymous:1949:ATAc

[Ano49c] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 28(3):596–598, July 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-3-596.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
3-596.pdf.

Anonymous:1949:ATAd

[Ano49d] Anonymous. Abstracts of technical articles by Bell System authors. The Bell System Technical Journal, 28(4):751–752, October 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-4-751.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
4-751.pdf.

Anonymous:1949:CIa

[Ano49e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 28(1):162–163, January 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/
                                        bstj28-1-162.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol28-1949/articles/bstj28-
                                        1-162.pdf.

Anonymous:1949:CIb

[Ano49f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 28(2):332–334, April 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/
                                        bstj28-2-332.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol28-1949/articles/bstj28-
                                        2-332.pdf.

Anonymous:1949:CIc

[Ano49g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 28(3):599–600, July 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/
                                        bstj28-3-599.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol28-1949/articles/bstj28-
                                        3-599.pdf.

Anonymous:1949:CId

[Ano49h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 28(4):753, October 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/

bstj28-4-753.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
4-753.pdf.

Anonymous:1949:ENR

[Ano49i] Anonymous. Editorial note regarding semiconductors. The Bell System Technical Journal, 28(3):335–343, July 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-3-335.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
3-335.pdf.

Armstrong:1946:XRS

[Arm46] Elizabeth J. Armstrong. X-ray studies of surface layers of crystals. The Bell System Technical Journal, 25(1):136–155, January 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-1-136.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
1-136.pdf.

Bond:1943:UXR

[BA43] W. L. Bond and E. J. Armstrong. Use of X-rays for determining the orientation of quartz crystals. The Bell System Technical Journal, 22(3):293–337, October 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/

                                        bstj22-3-293.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol22-1943/articles/bstj22-
                                        3-293.pdf.

Baldwin:1940:SSS

[Bal40] Millard W. Baldwin, Jr. The subjective sharpness of simulated television images. The Bell System Technical Journal, 19(4): 563–586, October 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol19/
                                        bstj19-4-563.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol19-1940/articles/bstj19-
                                        4-563.pdf.

Bardeen:1949:TCI

[Bar49] J. Bardeen. On the theory of the A-C impedance of a contact rectifier. The Bell System Technical Journal, 28(3):428–434, July 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/
                                        bstj28-3-428.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol28-1949/articles/bstj28-
                                        3-428.pdf.

Bardeen:1949:PPI

[BB49] J. Bardeen and W. H. Brattain. Physical principles involved in transistor action. The Bell System Technical Journal, 28 (2):239–277, April 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol28/
bstj28-2-239.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
2-239.pdf.

Bennett:1940:CMR

[Ben40] W. R. Bennett. Cross-modulation requirements on multichannel amplifiers below overload. The Bell System Technical Journal, 19(4): 587–610, October 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-4-587.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
4-587.pdf.

Bennett:1941:TDM

[Ben41] W. R. Bennett. Time division multiplex systems. The Bell System Technical Journal, 20 (2):199–221, April 1941. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-2-199.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
2-199.pdf.

Bennett:1944:RLR

[Ben44] W. R. Bennett. Response of a linear rectifier to signal and noise. The Bell System Technical Journal, 23(1):97–113, January 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol23/
                                        bstj23-1-97.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol23-
                                        1944/articles/bstj23-1-97.pdf.

Bennett:1947:BIR

[Ben47] W. R. Bennett. The biased ideal rectifier. The Bell System Technical Journal, 26(1): 139–169, January 1947. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol26/
                                        bstj26-1-139.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol26-1947/articles/bstj26-
                                        1-139.pdf.

Bennett:1948:SQS

[Ben48] W. R. Bennett. Spectra of quantized signals. The Bell System Technical Journal, 27 (3):446–472, July 1948. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-3-446.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        3-446.pdf.

Becker:1947:PUT

[BGP47] J. A. Becker, C. B. Green, and G. L. Pearson. Properties and uses of thermistors–thermally sensitive resistors. The Bell System Technical Journal, 26(1): 170–212, January 1947. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol26/
bstj26-1-170.pdf;   http://
www.alcatel-lucent.com/bstj/
vol26-1947/articles/bstj26-
1-170.pdf.

Blanchard:1941:HER

[Bla41] Julian Blanchard. The history of electrical resonance. The Bell System Technical Journal, 20(4): 415–433, October 1941. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-4-415.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
4-415.pdf.

Blackman:1943:EFI

[Bla43] R. B. Blackman. Effect of feedback on impedance. The Bell System Technical Journal, 22(3): 269–277, October 1943. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-3-269.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
3-269.pdf.

Blount:1941:DON

[Blo41] H. Blount. Design and operation of new copper wire drawing plant: Part I — design and operation of high speed copper wire drawing machines. The Bell System Technical Journal, 20 (1):95–110, January 1941. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol20/
                                        bstj20-1-95.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol20-
                                        1941/articles/bstj20-1-95.pdf.

Booth:1940:CBC

[BO40] R. P. Booth and T. M. Odarenko. Crosstalk between coaxial conductors in cable. The Bell System Technical Journal, 19 (3):358–384, July 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol19/
                                        bstj19-3-358.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol19-1940/articles/bstj19-
                                        3-358.pdf.

Bode:1940:RBA

[Bod40] H. W. Bode. Relations between attenuation and phase in feedback amplifier design. The Bell System Technical Journal, 19(3):421–454, July 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol19/
                                        bstj19-3-421.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol19-1940/articles/bstj19-
                                        3-421.pdf.

Bond:1943:MSQ

[Bon43a] W. L. Bond. Methods for specifying quartz crystal orientation and their determination by optical means. The Bell System Technical Journal, 22(2):224–262, July 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-2-224.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
2-224.pdf.

Bond:1943:MSP

[Bon43b] W. L. Bond. A mineral survey for piezo-electric materials. The Bell System Technical Journal, 22(2):145–152, July 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-2-145.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
2-145.pdf.

Bond:1943:MPP

[Bon43c] Walter L. Bond. The mathematics of physical properties of crystals. The Bell System Technical Journal, 22(1):1–72, January 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol22-
1943/articles/bstj22-1-1.pdf.

Bozorth:1940:PBF

[Boz40] R. M. Bozorth. The physical basis of ferromagnetism. The Bell System Technical Journal, 19(1):1–39, January 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol19/
                                        bstj19-1-1.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol19-
                                        1940/articles/bstj19-1-1.pdf.

Bennett:1949:EFV

[BP49] W. R. Bennett and L. C. Peterson. The electrostatic field in vacuum tubes with arbitrarily spaced elements. The Bell System Technical Journal, 28(2):303–314, April 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/
                                        bstj28-2-303.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol28-1949/articles/bstj28-
                                        2-303.pdf.

Brogle:1949:DRE

[Bro49] A. P. Brogle, Jr. The design of reactive equilizers. The Bell System Technical Journal, 28(4): 716–750, October 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/
                                        bstj28-4-716.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol28-1949/articles/bstj28-
                                        4-716.pdf.

Buckley:1942:FTT

[Buc42] Oliver E. Buckley. The future of transoceanic telephony. The Bell System Technical Journal, 21(1):1–19, June 1942. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol21/
bstj21-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol21-
1942/articles/bstj21-1-1.pdf.

Budenbom:1948:APW

[Bud48] H. T. Budenbom. Analysis and performance of waveguide-hybrid rings for microwaves. The Bell System Technical Journal, 27 (3):473–486, July 1948. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-3-473.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
3-473.pdf.

Burns:1940:MQC

[Bur40] G. K. Burns. Manufacture of quartz crystal filters. The Bell System Technical Journal, 19(4): 516–532, October 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-4-516.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
4-516.pdf.

Chinn:1940:NSV

[CGM40] H. A. Chinn, D. K. Gannett, and R. M. Morris. A new standard volume indicator and reference level. The Bell System Technical Journal, 19(1):94–137, January 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

                                        labs.com/BSTJ/images/Vol19/
                                        bstj19-1-94.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol19-
                                        1940/articles/bstj19-1-94.pdf.

Clarke:1946:SNE

[Cla46] J. L. Clarke. Some novel expressions for the propagation constant of a uniform line. The Bell System Technical Journal, 25 (1):156–157, January 1946. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol25/
                                        bstj25-1-156.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol25-1946/articles/bstj25-
                                        1-156.pdf.

Clos:1948:ADB

[Clo48] Charles Clos. An aspect of the dialing behavior of subscribers and its effect on the trunk plant. The Bell System Technical Journal, 27(3):424–445, July 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-3-424.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        3-424.pdf.

Cowan:1941:ERP

[CML41] F. A. Cowan, R. G. McCurdy, and I. E. Lattimer. Engineering requirements for program transmission circuits. The Bell System Technical Journal, 20 (2):235–249, April 1941. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-2-235.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
2-235.pdf.

Curtis:1940:CPT

[Cur40] A. M. Curtis. Contact phenomena in telephone switching circuits. The Bell System Technical Journal, 19(1):40–62, January 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-1-40.pdf;   http://www.
alcatel-lucent.com/bstj/vol19-
1940/articles/bstj19-1-40.pdf.

Darrow:1940:AI

[Dar40a] Karl K. Darrow. Analysis of the ionosphere. The Bell System Technical Journal, 19(3):455–488, July 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-3-455.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
3-455.pdf.

Darrow:1940:NF

[Dar40b] Karl K. Darrow. Nuclear fission. The Bell System Technical Journal, 19(2):267–289, April 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-2-267.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol19-1940/articles/bstj19-
                                        2-267.pdf.

Darrow:1941:FAW

[Dar41] Karl K. Darrow. Forces and atoms: The world of the physicist. The Bell System Technical Journal, 20(3):340–358, July 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol20/
                                        bstj20-3-340.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol20-1941/articles/bstj20-
                                        3-340.pdf.

Darrow:1942:E

[Dar42] Karl K. Darrow. Entropy. The Bell System Technical Journal, 21(1):51–74, June 1942. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol21/
                                        bstj21-1-51.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol21-
                                        1942/articles/bstj21-1-51.pdf.

Darrow:1943:MCS

[Dar43a] Karl K. Darrow. Memorial to the classical statistics. The Bell System Technical Journal, 22(1): 108–135, January 1943. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol22/
                                        bstj22-1-108.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol22-1943/articles/bstj22-
                                        1-108.pdf.

Darrow:1943:NSM

[Dar43b] Karl K. Darrow. The new statistical mechanics. The Bell System Technical Journal, 22(3): 362–392, October 1943. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-3-362.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
3-362.pdf.

Dennis:1949:RTM

[DF49] F. R. Dennis and E. P. Felch.

Reactance tube modulation of phase shift oscillators. The Bell System Technical Journal, 28(4): 601–607, October 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-4-601.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
4-601.pdf.

DHeedene:1944:EMD

[D’H44] A. R. D’Heedene. Effects of manufacturing deviations on crystal units for filters. The Bell System Technical Journal, 23 (3):260–281, July 1944. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-3-260.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
3-260.pdf.

Davey:1948:FST

[DM48] J. R. Davey and A. L. Matte. Frequency shift telegraphy – radio and wire applications. The Bell System Technical Journal, 27 (2):265–304, April 1948. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-2-265.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        2-265.pdf.

Dodge:1941:SSD

[DR41] H. F. Dodge and H. G. Romig. Single sampling and double inspection tables. The Bell System Technical Journal, 20(1):1–61, January 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol20/
                                        bstj20-1-1.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol20-
                                        1941/articles/bstj20-1-1.pdf.

Dagnall:1949:DEE

[DR49] C. H. Dagnall and P. W. Rounds. Delay equalization of eight-kilocycle carrier program circuits. The Bell System Technical Journal, 28(2):181–195, April 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/
                                        bstj28-2-181.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol28-1949/articles/bstj28-
                                        2-181.pdf.

Dudley:1940:CNS

[Dud40] Homer Dudley. The carrier nature of speech. The Bell System Technical Journal, 19(4):495–515, October 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-4-495.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
4-495.pdf.

Edson:1945:IBO

[Eds45] W. A. Edson. Intermittent behavior in oscillators. The Bell System Technical Journal, 24 (1):1–22, January 1945. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol24/
bstj24-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol24-
1945/articles/bstj24-1-1.pdf.

Eggleston:1945:ERB

[Egg45] Richard C. Eggleston. Evaluating the relative bending strength of crossarms. The Bell System Technical Journal, 24(1): 23–45, January 1945. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol24/
bstj24-1-23.pdf;   http://www.
alcatel-lucent.com/bstj/vol24-
1945/articles/bstj24-1-23.pdf.

Englund:1944:DCP

[Eng44] Carl R. Englund. Dielectric constants and power factors at centimeter wave-lengths. The Bell System Technical Journal, 23(1): 114–129, January 1944. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-1-114.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
1-114.pdf.

Fair:1945:PCO

[Fai45] I. E. Fair. Piezoelectric crystals in oscillator circuits. The Bell System Technical Journal, 24 (2):161–216, April 1945. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol24/
bstj24-2-161.pdf;   http://
www.alcatel-lucent.com/bstj/
vol24-1945/articles/bstj24-
2-161.pdf.

Fay:1947:HVO

[Fay47] C. E. Fay. High-vacuum oxidecathode pulse modulator tubes. The Bell System Technical Journal, 26(4):818–836, October 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-4-818.pdf;   http://
www.alcatel-lucent.com/bstj/
vol26-1947/articles/bstj26-
4-818.pdf.

Feldman:1949:BWT

[FB49] C. B. Feldman and W. R. Bennett. Band width and transmission performance. The Bell System Technical Journal, 28(3):490–595, July 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-3-490.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
3-490.pdf.

Fisk:1946:MGC

[FHH46] J. B. Fisk, H. D. Hagstrum, and P. L. Hartman. The magnetron as a generator of centimeter waves. The Bell System Technical Journal, 25(2):167–348, April 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-2-167.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
2-167.pdf.

Farkas:1949:BPF

[FHS49] F. S. Farkas, F. J. Hallenbeck, and F. E. Stehlik. Band pass filter, band elimination filter and phase simulating network for carrier program systems. The Bell System Technical Journal, 28 (2):196–220, April 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-2-196.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
2-196.pdf.

Friis:1947:RA

[FL47] H. T. Friis and W. D. Lewis. Radar antennas. The Bell System Technical Journal, 26(2): 219–317, April 1947. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol26/
                                        bstj26-2-219.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol26-1947/articles/bstj26-
                                        2-219.pdf.

Ford:1946:CVT

[For46] G. T. Ford. Characteristics of vacuum tubes for radar intermediate frequency amplifiers. The Bell System Technical Journal, 25(3):385–407, July 1946. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol25/
                                        bstj25-3-385.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol25-1946/articles/bstj25-
                                        3-385.pdf.

Friis:1948:MRR

[Fri48] H. T. Friis. Microwave repeater research. The Bell System Technical Journal, 27(2):183–246, April 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-2-183.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        2-183.pdf.

Fry:1941:IM

[Fry41] Thornton C. Fry. Industrial mathematics. The Bell System Technical Journal, 20(3):255–292, July 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-3-255.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
3-255.pdf.

Fuller:1946:SRC

[Ful46] C. S. Fuller. Some recent contributions to synthetic rubber research. The Bell System Technical Journal, 25(3):351–384, July 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-3-351.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
3-351.pdf.

Green:1946:TFM

[GFF46] E. I. Green, H. J. Fisher, and J. G. Ferguson. Techniques and facilities for microwave radar testing. The Bell System Technical Journal, 25(3):435–482, July 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-3-435.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
3-435.pdf.

Goucher:1946:SGS

[GHDR46] F. S. Goucher, J. R. Haynes, W. A. Depp, and E. J. Ryder. Spark gap switches for radar. The Bell System Technical Journal, 25(4):563–602, October 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-4-563.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
4-563.pdf.

Gannett:1944:DCP

[GK44] D. K. Gannett and Iden Kerney. The discernibility of changes in program band width. The Bell System Technical Journal, 23(1):1–10, January 1944. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol23-
1944/articles/bstj23-1-1.pdf.

Goodall:1947:TPC

[Goo47] W. M. Goodall. Telephony by pulse code modulation. The Bell System Technical Journal, 26(3):395–409, July 1947. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-3-395.pdf;   http://
www.alcatel-lucent.com/bstj/
vol26-1947/articles/bstj26-
3-395.pdf.

Gould:1940:CCC

[Gou40] K. E. Gould. Crosstalk in coaxial cables — analysis based on short-circuited and open tertiaries. The Bell System Technical Journal, 19(3):341–357, July 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-3-341.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
3-341.pdf.

Graham:1946:LST

[Gra46] Robert E. Graham. Linear servo theory. The Bell System Technical Journal, 25(4):616–651, October 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-4-616.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
4-616.pdf.

Greenidge:1944:MFP

[Gre44] R. M. C. Greenidge. The mounting and fabrication of plated quartz crystal units. The Bell System Technical Journal, 23 (3):234–259, July 1944. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-3-234.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
3-234.pdf.

Germer:1940:TDI

[GS40] L. H. Germer and K. H. Storks. Technical digests: An interesting application of electron diffraction. The Bell System Technical Journal, 19(1):152–155, January 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-1-152.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
1-152.pdf.

Gray:1948:ASL

[GS48] Marion C. Gray and S. A. Schelkunoff. The approximate solution of linear differential equations. The Bell System Technical Journal, 27(2):350–364, April 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-2-350.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
2-350.pdf.

Hartley:1941:SSD

[Har41] R. V. L. Hartley. Steady state delay as related to aperiodic signals. The Bell System Technical Journal, 20(2):222–234, April 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-2-222.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
2-222.pdf.

Harrison:1945:MPI

[Har45] C. W. Harrison. The measurement of the performance index of quartz plates. The Bell System Technical Journal, 24(2):217–252, April 1945. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol24/
bstj24-2-217.pdf;   http://
www.alcatel-lucent.com/bstj/
vol24-1945/articles/bstj24-
2-217.pdf.

Hebbert:1941:TCT

[Heb41] C. M. Hebbert. The transmission characteristics of toll telephone cables at carrier frequencies. The Bell System Technical Journal, 20(3):293–330, July 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-3-293.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
3-293.pdf.

Heising:1940:REL

[Hei40] R. A. Heising. Radio extension links to the telephone system. The Bell System Technical Journal, 19(4):611–646, October 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-4-611.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
4-611.pdf.

Herring:1949:TTP

[Her49] Conyers Herring. Theory of transient phenomena in the transport of holes in an excess semiconductor. The Bell System Technical Journal, 28(3):401–427, July 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-3-401.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
3-401.pdf.

Hollenberg:1949:EOA

[Hol49] A. V. Hollenberg. Experimental observation of amplification by interaction between two electron streams. The Bell System Technical Journal, 28(1):52–58, January 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-1-52.pdf;   http://www.
alcatel-lucent.com/bstj/vol28-
1949/articles/bstj28-1-52.pdf.

Hoyt:1947:PFM

[Hoy47] Ray S. Hoyt. Probability functions for the modulus and angle of the normal complex variate. The Bell System Technical Journal, 26(2):318–359, April 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-2-318.pdf;   http://
www.alcatel-lucent.com/bstj/
vol26-1947/articles/bstj26-
2-318.pdf.

Johnson:1946:DT

[Joh46] K. S. Johnson. Decibel tables. The Bell System Technical Journal, 25(1):158, January 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-1-158.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
1-158.pdf.

Jones:1946:PCV

[JP46] T. A. Jones and K. W. Pfleger.

Performance characteristics of various carrier telegraph methods. The Bell System Technical Journal, 25(3):483–531, July 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-3-483.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
3-483.pdf.

Keeling:1942:UDS

[KC42] D. B. Keeling and L. E. Cisne.

Using double sampling inspection in a manufacturing plant. The Bell System Technical Journal, 21(1):37–50, June 1942. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol21/
bstj21-1-37.pdf;   http://www.
alcatel-lucent.com/bstj/vol21-
1942/articles/bstj21-1-37.pdf.

Kock:1948:MDL

[Koc48] Winston E. Kock. Metallic delay lenses. The Bell System Technical Journal, 27(1):58–82, January 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-1-58.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol27-
                                        1948/articles/bstj27-1-58.pdf.

Kinzer:1947:EPS

[KW47a] J. P. Kinzer and I. G. Wilson. End plate and side wall currents in circular cylinder cavity resonator. The Bell System Technical Journal, 26(1):31–79 (1 plate), January 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol26/
                                        bstj26-1-31.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol26-
                                        1947/articles/bstj26-1-31.pdf.

Kinzer:1947:SRC

[KW47b] J. P. Kinzer and I. G. Wilson. Some results on cylindrical cavity resonators. The Bell System Technical Journal, 26(3):410– 445 (1 plate), July 1947. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol26/
                                        bstj26-3-410.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol26-1947/articles/bstj26-
                                        3-410.pdf.

LeCorbeiller:1943:EWR

[LeC43] P. LeCorbeiller. Electromagnetic waves: Review of S. A. Schelkunoff’s book. The Bell System Technical Journal, 22(3): 393–396, October 1943. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-3-393.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
3-393.pdf.

Leutritz:1946:WSC

[Leu46] John Leutritz, Jr. A wood soil contact culture technique for laboratory study of wooddestroying fungi, wood decay and wood preservation. The Bell System Technical Journal, 25(1): 102–135, January 1946. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-1-102.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
1-102.pdf.

Legg:1940:CPM

[LG40] V. E. Legg and F. J. Given.

Compressed powered molybdenum permalloy for high quality inductance coils. The Bell System Technical Journal, 19(3):385–406, July 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-3-385.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol19-1940/articles/bstj19-
                                        3-385.pdf.

Little:1941:ITW

[Lit41] J. S. Little. Insulation of telephone wire with paper pulp. The Bell System Technical Journal, 20 (1):82–94, January 1941. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol20/
                                        bstj20-1-82.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol20-
                                        1941/articles/bstj20-1-82.pdf.

Lozier:1947:SAP

[Loz47] J. C. Lozier. Spectrum analysis of pulse modulated waves. The Bell System Technical Journal, 26(2):360–387, April 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol26/
                                        bstj26-2-360.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol26-1947/articles/bstj26-
                                        2-360.pdf.

Leconte:1949:CSC

[LPSW49] R. A. Leconte, D. B. Penick, C. W. Schramm, and A. J. Wier. A carrier system for 8000-cycle program transmission. The Bell System Technical Journal, 28 (2):165–180, April 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/
                                        bstj28-2-165.pdf;   http://

www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
2-165.pdf.

Lewis:1948:NRB

[LT48] W. D. Lewis and L. C. Tillotson. A non-reflecting branching filter for microwaves. The Bell System Technical Journal, 27 (1):83–95, January 1948. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-1-83.pdf;   http://www.
alcatel-lucent.com/bstj/vol27-
1948/articles/bstj27-1-83.pdf.

MacColl:1943:FEE

[Mac43] L. A. MacColl. The fundamental equations of electron motion. Dynamics of high speed particles. The Bell System Technical Journal, 22(2):153–177, July 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-2-153.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
2-153.pdf.

Marrison:1948:EQC

[Mar48] Warren A. Marrison. The evolution of the quartz crystal clock. The Bell System Technical Journal, 27(3):510–588, July 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-3-510.pdf;   http://

                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        3-510.pdf.

Mason:1940:LTC

[Mas40] W. P. Mason. Low temperature coefficient quartz crystals. The Bell System Technical Journal, 19(1):74–93, January 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol19/
                                        bstj19-1-74.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol19-
                                        1940/articles/bstj19-1-74.pdf.

Mason:1941:EMA

[Mas41] W. P. Mason. Electrical and mechanical analogies. The Bell System Technical Journal, 20(4): 405–414, October 1941. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol20/
                                        bstj20-4-405.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol20-1941/articles/bstj20-
                                        4-405.pdf.

Mason:1943:QCA

[Mas43] W. P. Mason. Quartz crystal applications. The Bell System Technical Journal, 22(2):178–223, July 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol22/
                                        bstj22-2-178.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol22-1943/articles/bstj22-
                                        2-178.pdf.

Mason:1947:FSO

[Mas47] W. P. Mason. First and second order equations for piezoelectric crystals expressed in tensor form. The Bell System Technical Journal, 26(1):80–138, January 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-1-80.pdf;   http://www.
alcatel-lucent.com/bstj/vol26-
1947/articles/bstj26-1-80.pdf.

Matte:1940:ACT

[Mat40] A. L. Matte. Advances in carrier telegraph transmission. The Bell System Technical Journal, 19 (2):161–208, April 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-2-161.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
2-161.pdf.

McSkimin:1944:TAM

[McS44] H. J. McSkimin. Theoretical analysis of modes of vibration for isotropic rectangular plates having all surfaces free. The Bell System Technical Journal, 23 (2):151–177, April 1944. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-2-151.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
2-151.pdf.

Mindlin:1945:DPC

[Min45] Raymond D. Mindlin. Dynamics of package cushioning. The Bell System Technical Journal, 24 (3):353–461, July/October 1945. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol24/
                                        bstj24-3-353.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol24-1945/articles/bstj24-
                                        3-353.pdf.

Morrison:1947:RR

[Mor47] L. W. Morrison, Jr. The radar receiver. The Bell System Technical Journal, 26(4):693–817, October 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol26/
                                        bstj26-4-693.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol26-1947/articles/bstj26-
                                        4-693.pdf.

Mott:1944:IRT

[Mot44] E. E. Mott. Indicial response of telephone receivers. The Bell System Technical Journal, 23 (2):135–150, April 1944. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol23/
                                        bstj23-2-135.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol23-1944/articles/bstj23-
                                        2-135.pdf.

Meacham:1948:EMP

[MP48] L. A. Meacham and E. Peterson.

An experimental multichannel pulse code modulation system of toll quality. The Bell System Technical Journal, 27(1):1–43, January 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol27-
1948/articles/bstj27-1-1.pdf.

Mason:1940:EWF

[MS40] W. P. Mason and R. A. Sykes.

Electrical wave filters employing crystals with normal and divided electrodes. The Bell System Technical Journal, 19 (2):221–248, April 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-2-221.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
2-221.pdf.

Mueller:1947:PA

[MT47] G. E. Mueller and W. A. Tyrrell.

Polyrod antennas. The Bell System Technical Journal, 26(4): 837–851, October 1947. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-4-837.pdf;   http://
www.alcatel-lucent.com/bstj/
vol26-1947/articles/bstj26-
4-837.pdf.

Muller:1948:TMS

[Mul48] J. T. Muller. Transients in mechanical systems. The Bell System Technical Journal, 27(4): 657–683, October 1948. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-4-657.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        4-657.pdf.

Mumford:1948:MFF

[Mum48] W. W. Mumford. Maximallyflat filters in waveguide. The Bell System Technical Journal, 27 (4):684–713, October 1948. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-4-684.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        4-684.pdf.

Mumford:1949:BBM

[Mum49] W. W. Mumford. A broad-band microwave noise source. The Bell System Technical Journal, 28(4): 608–618, October 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/
                                        bstj28-4-608.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol28-1949/articles/bstj28-
                                        4-608.pdf.

Nyquist:1940:EQC

[NP40] H. Nyquist and K. W. Pfleger.

Effect of the quadrature component in single sideband transmission. The Bell System Technical Journal, 19(1):63–73, January 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-1-63.pdf;   http://www.
alcatel-lucent.com/bstj/vol19-
1940/articles/bstj19-1-63.pdf.

Padowicz:1942:DDH

[Pad42] H. N. Padowicz. Diamond dies for high-speed drawing of copper wire. The Bell System Technical Journal, 21(1):20–36, June 1942. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol21/
bstj21-1-20.pdf;   http://www.
alcatel-lucent.com/bstj/vol21-
1942/articles/bstj21-1-20.pdf.

Peterson:1946:CPR

[Pet46] E. Peterson. Coil pulsers for radar. The Bell System Technical Journal, 25(4):603–615, October 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-4-603.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
4-603.pdf.

Peterson:1948:ECL

[Pet48] Liss C. Peterson. Equivalent circuits of linear active four-terminal networks. The Bell System Technical Journal, 27(4):593–622, October 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-4-593.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
4-593.pdf.

Pierce:1949:NTH

[PH49] J. R. Pierce and W. B. Hebenstreit. A new type of highfrequency amplifier. The Bell System Technical Journal, 28 (1):33–51, January 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-1-33.pdf;   http://www.
alcatel-lucent.com/bstj/vol28-
1949/articles/bstj28-1-33.pdf.

Pierce:1943:NTL

[Pie43] J. R. Pierce. A note on the transmission line equation in terms of impedance. The Bell System Technical Journal, 22(2):263–265, July 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-2-263.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
2-263.pdf.

Pierce:1945:PLE

[Pie45] J. R. Pierce. Physical limitations in electron ballistics. The Bell System Technical Journal, 24 (3):305–321, July/October 1945. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol24/
bstj24-3-305.pdf;   http://
www.alcatel-lucent.com/bstj/
vol24-1945/articles/bstj24-
3-305.pdf.

Pierce:1948:NRE

[Pie48a] J. R. Pierce. Noise in resistances and electron streams. The Bell System Technical Journal, 27 (1):158–174, January 1948. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-1-158.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
1-158.pdf.

Pierce:1948:TFT

[Pie48b] J. R. Pierce. Transverse fields in traveling-wave tubes. The Bell System Technical Journal, 27(4): 732–746, October 1948. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-4-732.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
4-732.pdf.

Polkinghorn:1940:SSM

[Pol40] F. A. Polkinghorn. A single sideband musa receiving system for commercial operation on transatlantic radio telephone circuits.

The Bell System Technical Journal, 19(2):306–335, April 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol19/
                                        bstj19-2-306.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol19-1940/articles/bstj19-
                                        2-306.pdf.

Pomeroy:1947:PMI

[Pom47] Allen F. Pomeroy. Precision measurement of impedance mismatches in waveguide. The Bell System Technical Journal, 26(3):446–459, July 1947. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol26/
                                        bstj26-3-446.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol26-1947/articles/bstj26-
                                        3-446.pdf.

Pierce:1947:RO

[PS47] J. R. Pierce and W. G. Shepherd. Reflex oscillators. The Bell System Technical Journal, 26(3):460–681, July 1947. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol26/
                                        bstj26-3-460.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol26-1947/articles/bstj26-
                                        3-460.pdf.

Rappleye:1946:SDE

[Rap46] S. C. Rappleye. A study of the delays encountered by toll operators in obtaining an idle trunk. The Bell System Technical Journal, 25(4):539–562, October 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-4-539.pdf;   http://
www.alcatel-lucent.com/bstj/
vol25-1946/articles/bstj25-
4-539.pdf.

Rea:1944:ETD

[Rea44] W. T. Rea. Effect of telegraph distortion on the margins of operation of start-stop receivers. The Bell System Technical Journal, 23(3):207–233, July 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-3-207.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
3-207.pdf.

Rice:1941:SSS

[Ric41] S. O. Rice. Steady state solutions of transmission line equations. The Bell System Technical Journal, 20(2):131–178, April 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-2-131.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
2-131.pdf.

Rice:1944:MAR

[Ric44] S. O. Rice. Mathematical analysis of random noise. The Bell System Technical Journal, 23 (3):282–332, July 1944. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol23/
                                        bstj23-3-282.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol23-1944/articles/bstj23-
                                        3-282.pdf.

Rice:1945:MAR

[Ric45] S. O. Rice. Mathematical analysis of random noise. The Bell System Technical Journal, 24(1): 46–156, January 1945. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol24/
                                        bstj24-1-46.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol24-
                                        1945/articles/bstj24-1-46.pdf.

Rice:1948:RCB

[Ric48a] S. O. Rice. Reflections from circular bends in rectangular wave guides. Matrix theory. The Bell System Technical Journal, 27 (2):305–349, April 1948. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-2-305.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        2-305.pdf.

Rice:1948:SPS

[Ric48b] S. O. Rice. Statistical properties of a sine wave plus random noise. The Bell System Technical Journal, 27(1):109–157, January 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-1-109.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
1-109.pdf.

Rice:1949:RCR

[Ric49a] S. O. Rice. Reflection from corners in rectangular wave guides— conformal transformation. The Bell System Technical Journal, 28 (1):104–135, January 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-1-104.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
1-104.pdf.

Rice:1949:SSO

[Ric49b] S. O. Rice. A set of secondorder differential equations associated with reflections in rectangular wave guides application to guide connected to horn. The Bell System Technical Journal, 28 (1):136–156, January 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-1-136.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
1-136.pdf.

Rideout:1948:NPT

[Rid48] V. C. Rideout. A note on paralleltuned transformer design. The Bell System Technical Journal, 27 (1):96–108, January 1948. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-1-96.pdf;   http://www.
                                        alcatel-lucent.com/bstj/vol27-
                                        1948/articles/bstj27-1-96.pdf.

Ring:1948:MDD

[Rin48] D. H. Ring. The measurement of delay distortion in microwave repeaters. The Bell System Technical Journal, 27(2):247–264, April 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-2-247.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        2-247.pdf.

Ryder:1949:SCA

[RK49] R. M. Ryder and R. J. Kircher. Some circuit aspects of the transistor. The Bell System Technical Journal, 28(3):367–400, July 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol28/
                                        bstj28-3-367.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol28-1949/articles/bstj28-
                                        3-367.pdf.

Robertson:1947:TMR

[Rob47] Sloan D. Robertson. Targets for microwave radar navigation. The Bell System Technical Journal, 26(4):852–869, October 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-4-852.pdf;   http://
www.alcatel-lucent.com/bstj/
vol26-1947/articles/bstj26-
4-852.pdf.

Robertson:1949:EAP

[Rob49a] Sloan D. Robertson. Electronic admittances of parallel-plane electron tubes at 4000 megacycles. The Bell System Technical Journal, 28(4):619–646, October 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-4-619.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
4-619.pdf.

Robertson:1949:MMP

[Rob49b] Sloan D. Robertson. A method of measuring phase at microwave frequencies. The Bell System Technical Journal, 28(1):99–103, January 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-1-99.pdf;   http://www.
alcatel-lucent.com/bstj/vol28-
1949/articles/bstj28-1-99.pdf.

Robertson:1949:PFP

[Rob49c] Sloan D. Robertson. Passive fourpole admittances of microwave triodes. The Bell System Technical Journal, 28(4):647–655, October 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-4-647.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
4-647.pdf.

Samuel:1945:EBH

[Sam45] A. L. Samuel. Electron ballistics in high-frequency fields. The Bell System Technical Journal, 24(3):322–352, July/October 1945. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol24/
bstj24-3-322.pdf;   http://
www.alcatel-lucent.com/bstj/
vol24-1945/articles/bstj24-
3-322.pdf.

Schumacher:1941:RVT

[SB41] Earle E. Schumacher and G. M. Bouton. A rapid visual test for the quantitative determination of small concentrations of calcium in lead. The Bell System Technical Journal, 20(4):434–438, October 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-4-434.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
4-434.pdf.

Schumacher:1943:MFW

[SBP43] E. E. Schumacher, G. M. Bouton, and G. S. Phipps. The metallurgy of fillet wiped soldered joints. The Bell System Technical Journal, 22(1):73–79, January 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-1-73.pdf;   http://www.
alcatel-lucent.com/bstj/vol22-
1943/articles/bstj22-1-73.pdf.

Schelkunoff:1943:MTL

[Sch43] S. A. Schelkunoff. A mathematical theory of linear arrays. The Bell System Technical Journal, 22(1):80–107, January 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-1-80.pdf;   http://www.
alcatel-lucent.com/bstj/vol22-
1943/articles/bstj22-1-80.pdf.

Schelkunoff:1948:MEF

[Sch48]  S. A. Schelkunoff.  Methods

of electromagnetic field analysis. The Bell System Technical Journal, 27(3):487–509, July 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-3-487.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
3-487.pdf.

Samuel:1946:GDT

[SCM46] A. L. Samuel, J. W. Clark, and W. W. Mumford. The gasdischarge transmit-receive switch. The Bell System Technical Journal, 25(1):48–101, January 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-1-48.pdf;   http://www.
alcatel-lucent.com/bstj/vol25-
1946/articles/bstj25-1-48.pdf.

Schumacher:1940:MMT

[SE40] Earle E. Schumacher and W. C. Ellis. Metallic materials in the telephone system. The Bell System Technical Journal, 19(1): 138–151, January 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-1-138.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
1-138.pdf.

Sears:1948:EBD

[Sea48] R. W. Sears. Electron beam deflection tube for pulse code modulation. The Bell System Technical Journal, 27(1):44–57, January 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-1-44.pdf;   http://www.
alcatel-lucent.com/bstj/vol27-
1948/articles/bstj27-1-44.pdf.

Shaw:1944:CDW

[Sha44] Thomas Shaw. The conquest of distance by wire telephony. The Bell System Technical Journal, 23(4):337–421, October 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-4-337.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
4-337.pdf.

Shannon:1945:MTC

[Sha45] Claude Shannon. A mathematical theory of cryptography. Classified report, Bell Laboratories, Murray Hill, NJ, USA, September 1, 1945.

Shannon:1948:MTCa

[Sha48a] C. E. Shannon. A mathematical theory of communication. The Bell System Technical Journal, 27(3):379–423, July 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-3-379.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
3-379.pdf.

Shannon:1948:MTCb

[Sha48b] C. E. Shannon. A mathematical theory of communication. The Bell System Technical Journal, 27(4):623–656, October 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-

labs.com/BSTJ/images/Vol27/
bstj27-4-623.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
4-623.pdf.

Shannon:1949:CTS

[Sha49a] Claude E. Shannon. Communication theory of secrecy systems. The Bell System Technical Journal, 28(4):656–715, October 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://en.

wikipedia.org/wiki/Communication_
Theory_of_Secrecy_Systems;
http://www.alcatel-lucent.com/
bstj/vol28-1949/articles/bstj28-
4-656.pdf;   http://www.cs.
ucla.edu/~jkong/research/security/
shannon1949.pdf. A footnote on

the initial page says: “The material in this paper appeared in a confidential report, ‘A Mathematical Theory of Cryptography’, dated Sept. 1, 1945 ([Sha45]), which has now been declassified.”.

Shannon:1949:STT

[Sha49b] Claude E. Shannon. The synthesis of two terminal switching circuits. The Bell System Technical Journal, 28(1):59–98, January 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-1-59.pdf;   http://www.
alcatel-lucent.com/bstj/vol28-
1949/articles/bstj28-1-59.pdf.

Shockley:1949:TPJ

[Sho49] W. Shockley. The theory of p–n junctions in semiconductors and p–n junction transistors. The Bell System Technical Journal, 28(3):435–489, July 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-3-435.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
3-435.pdf.

Skellett:1940:CIO

[Ske40] A. M. Skellett. The Coronaviser, an instrument for observing the solar corona in full sunlight. The Bell System Technical Journal, 19(2):249–261, April 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-2-249.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
2-249.pdf.

Skellett:1944:MFR

[Ske44] A. M. Skellett. The magnetically focused radial beam vacuum tube. The Bell System Technical Journal, 23(2):190–202, April 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-2-190.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
2-190.pdf.

Slonczewski:1940:HAH

[Slo40]  T. Slonczewski.  High accu-

racy heterodyne oscillators. The Bell System Technical Journal, 19(3):407–420, July 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-3-407.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
3-407.pdf.

Slonczewski:1949:TCE

[Slo49] T. Slonczewski. Transconductance as a criterion of electron tube performance. The Bell System Technical Journal, 28 (2):315–328, April 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-2-315.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
2-315.pdf.

Steinberg:1940:RWF

[SMG40] J. C. Steinberg, H. C. Montgomery, and M. B. Gardner. Results of the World’s Fair hearing tests. The Bell System Technical Journal, 19(4):533–562, October 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-4-533.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
4-533.pdf.

Scaff:1947:DSC

[SO47] J. H. Scaff and R. S. Ohl. Development of silicon crystal rectifiers for microwave radar receivers. The Bell System Technical Journal, 26(1):1–30, January 1947. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol26-
1947/articles/bstj26-1-1.pdf.

Schumacher:1940:LTA

[SP40] Earle E. Schumacher and G. S.

Phipps. Lead–tin–arsenic wiping solder. The Bell System Technical Journal, 19(2):262–266, April 1940. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-2-262.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
2-262.pdf.

Shockley:1949:HIG

[SPH49] W. Shockley, G. L. Pearson, and J. R. Haynes. Hole injection in germanium — quantitative studies and filamentary transistors. The Bell System Technical Journal, 28(3):344–366, July 1949. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-3-344.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
3-344.pdf.

Schumacher:1944:SAP

[SS44] Earle E. Schumacher and Alexander G. Souden. Some aspects of powder metallurgy. The Bell System Technical Journal, 23(4): 422–457, October 1944. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol23/
                                        bstj23-4-422.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol23-1944/articles/bstj23-
                                        4-422.pdf.

Stansel:1941:SAW

[Sta41] F. R. Stansel. Some analyses of wave shapes used in harmonic producers. The Bell System Technical Journal, 20(3):331–339, July 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol20/
                                        bstj20-3-331.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol20-1941/articles/bstj20-
                                        3-331.pdf.

Sunde:1945:LPB

[Sun45] E. D. Sunde. Lightning protection of buried toll cable. The Bell System Technical Journal, 24 (2):253–300, April 1945. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol24/
                                        bstj24-2-253.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol24-1945/articles/bstj24-
                                        2-253.pdf.

Strieby:1941:TTW

[SW41] M. E. Strieby and J. F. Wentz.

Television transmission over wire lines. The Bell System Technical Journal, 20(1):62–81, January 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-1-62.pdf;   http://www.
alcatel-lucent.com/bstj/vol20-
1941/articles/bstj20-1-62.pdf.

Sykes:1944:MMQ

[Syk44a] R. A. Sykes. Modes of motion in quartz crystals, the effects of coupling and methods of design. The Bell System Technical Journal, 23(1):52–96, January 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-1-52.pdf;   http://www.
alcatel-lucent.com/bstj/vol23-
1944/articles/bstj23-1-52.pdf.

Sykes:1944:PMQ

[Syk44b] R. A. Sykes. Principles of mounting quartz plates. The Bell System Technical Journal, 23 (2):178–189, April 1944. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-2-178.pdf;   http://
www.alcatel-lucent.com/bstj/
vol23-1944/articles/bstj23-
2-178.pdf.

Tinus:1946:EFC

[TH46] W. C. Tinus and W. H. C. Higgins. Early fire-control radars for naval vessels. The Bell System Technical Journal, 25(1):1–47, January 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol25/
bstj25-1-1.pdf;   http://www.
alcatel-lucent.com/bstj/vol25-
1946/articles/bstj25-1-1.pdf.

Thomas:1947:TPA

[Tho47] D. E. Thomas. Tables of phase associated with a semi-infinite unit slope of attenuation. The Bell System Technical Journal, 26(4): 870–899, October 1947. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol26/
bstj26-4-870.pdf;   http://
www.alcatel-lucent.com/bstj/
vol26-1947/articles/bstj26-
4-870.pdf.

Trueblood:1949:LCO

[TS49] H. M. Trueblood and E. D. Sunde. Lightning current observations in buried cable. The Bell System Technical Journal, 28 (2):278–302, April 1949. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol28/
bstj28-2-278.pdf;   http://
www.alcatel-lucent.com/bstj/
vol28-1949/articles/bstj28-
2-278.pdf.

Vaage:1940:SFT

[Vaa40] E. F. Vaage. A solution for faults at two locations in threephase power systems. The Bell System Technical Journal, 19 (2):290–305, April 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-2-290.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
2-290.pdf.

VanWynen:1943:DTT

[Van43] K. G. VanWynen. Design of two-terminal balancing networks. The Bell System Technical Journal, 22(3):278–292, October 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-3-278.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
3-278.pdf.

Wade:1940:EDT

[Wad40] L. G. Wade. Electrical drying of telephone cable. The Bell System Technical Journal, 19 (2):209–220, April 1940. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol19/
bstj19-2-209.pdf;   http://
www.alcatel-lucent.com/bstj/
vol19-1940/articles/bstj19-
2-209.pdf.

Werring:1941:EPD

[Wer41] W. W. Werring. Engineering problems in dimensions and tolerances. The Bell System Technical Journal, 20(2):179–198, April 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-2-179.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
2-179.pdf.

Wilkinson:1941:RHT

[Wil41a] Roger I. Wilkinson. The reliability of holding time measurements. The Bell System Technical Journal, 20(4):365–404, October 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-4-365.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
4-365.pdf.

Wiltrakis:1941:DON

[Wil41b] J. E. Wiltrakis. Design and operation of new copper wire drawing plant: Part II — equipping and operating the new wire mill. The Bell System Technical Journal, 20(1):111–124, January 1941. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol20/
bstj20-1-111.pdf;   http://
www.alcatel-lucent.com/bstj/
vol20-1941/articles/bstj20-
1-111.pdf.

Willard:1943:RQI

[Wil43] G. W. Willard. Raw quartz, its imperfections and inspection. The Bell System Technical Journal, 22(3):338–361, October 1943. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol22/
bstj22-3-338.pdf;   http://
www.alcatel-lucent.com/bstj/
vol22-1943/articles/bstj22-
3-338.pdf.

Willard:1944:UET

[Wil44] G. W. Willard. Use of the etch technique for determining orientation and twinning in quartz crystals. The Bell System Technical Journal, 23(1):11–51, January 1944. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol23/
bstj23-1-11.pdf;   http://www.
alcatel-lucent.com/bstj/vol23-
1944/articles/bstj23-1-11.pdf.

Wise:1948:PCG

[Wis48] W. Howard Wise. Potential coefficients for ground return circuits. The Bell System Technical Journal, 27(2):365–371, April 1948. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.bell-
labs.com/BSTJ/images/Vol27/
bstj27-2-365.pdf;   http://
www.alcatel-lucent.com/bstj/
vol27-1948/articles/bstj27-
2-365.pdf.

Wilson:1946:HQR

[WSK46] I. G. Wilson, C. W. Schramm, and J. P. Kinzer. High Q resonant cavities for microwave testing. The Bell System Technical Journal, 25(3):408–434, July 1946. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol25/
                                        bstj25-3-408.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol25-1946/articles/bstj25-
                                        3-408.pdf.

Zinn:1948:TRF

[Zin48] Manvel K. Zinn. Transient response of an FM receiver. The Bell System Technical Journal, 27 (4):714–731, October 1948. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.bell-
                                        labs.com/BSTJ/images/Vol27/
                                        bstj27-4-714.pdf;   http://
                                        www.alcatel-lucent.com/bstj/
                                        vol27-1948/articles/bstj27-
                                        4-714.pdf.